# DiTCtrl: Exploring Attention Control in Multi-Modal Diffusion Transformer for Tuning-Free Multi-Prompt Longer Video Generation

### v1

@@ Abstract
Sora-like video generation models have achieved remarkable progress with a Multi-Modal Diffusion Transformer (MM-DiT) architecture. However, the current video generation models predominantly focus on single-prompt, struggling to generate coherent scenes with multiple sequential prompts that better reflect real-world dynamic scenarios. While some pioneering works have explored multi-prompt video generation, they face significant challenges including strict training data requirements, weak prompt following, and unnatural transitions. To address these problems, we propose DiTCtrl, a training-free multi-prompt video generation method under MM-DiT architectures for the first time. Our key idea is to take the multi-prompt video generation task as temporal video editing with smooth transitions. To achieve this goal, we first analyze MM-DiT’s attention mechanism, finding that the 3D full attention behaves similarly to that of the cross/self-attention blocks in the UNet-like diffusion models, enabling mask-guided precise semantic control across different prompts with attention sharing for multi-prompt video generation. Based on our careful design, the video generated by DiTCtrl achieves smooth transitions and consistent object motion given multiple sequential prompts without additional training. Besides, we also present MPVBench, a new benchmark specially designed for multi-prompt video generation to evaluate the performance of multi-prompt generation. Extensive experiments demonstrate that our method achieves state-of-the-art performance without additional training.
Figure 1: The proposed DiTCtrl takes multiple text prompts as input and demonstrates superior capability in generating longer videos with complex motion and smooth transitions as output. In this figure, we showcase a challenging example where an athlete glides through three distinct scenes. Despite the complex subject motion and dramatic camera movement, our method maintains remarkable stability throughout the sequence and seamless semantic transitions that faithfully follow the prompt descriptions.
@@ 1 Introduction
Text-to-video (T2V) generation has made remarkable progress in the AIGC era [10, 47, 41], and breakthroughs such as Sora [28] have demonstrated impressive capabilities in generating longer videos through DiT [30] architecture and large-scale per-taining. However, feeding sequential prompts into current state-of-the-art video generation models (e.g., Kling [3], Gen3 [2], CogVideoX [45]) fails to produce coherent video sequences that exhibit natural transitions with precise prompt-following. This limitation stems from their fundamental design and single-prompt training paradigm, making them inadequate for depicting real-world scenarios’ dynamic, multi-action nature.
@@
  • We pioneers the analysis of MM-DiT’s attention mechanism, finding that 3D full atteniton has similar behaviors to that of the cross/self-attention blocks in the UNet-like diffusion models, enabling mask-guided precise semantic control across different prompts for enhanced generation consistency.
@@ 2 Related Work
Training diffusion models on long videos often demand significant computational resources. Consequently, current video diffusion models are typically trained on videos with a limited number of frames. As a result, the quality of generated videos often degrades significantly during inference when generating longer videos. To address this problem, some works [18, 43, 39, 17] employ an autoregressive mechanism for long video generation. However, due to error accumulation, these methods often suffer from quality degradation after a few iterations. Alternatively, tuning-free methods [40, 33, 38, 21, 5] have been developed to extend off-the-shelf short-video diffusion models for generating long videos without additional training. For instance, Gen-L-Video [40] processes long videos as short video clips with temporal overlapping during the denoising process. FreeNoise [33] explores the influence of initial noises and conducts temporal attention fusion based on the sliding window for temporal consistency. Inspired by these works, we propose a novel KV-sharing mechanism and latent blending strategy for seamless transitions between different segments without additional training.
@@
Attention control is gaining popularity due to its ability to perform zero-shot image or video editing without the need for additional data. In the realm of image editing, MasaCtrl [9] enhances the existing self-attention mechanism in diffusion models by introducing mutual self-attention. This allows for querying correlated content and textures from source images, ensuring consistent and coherent edits. Prompt-to-Prompt [19] utilizes cross-attention layers to control the relation between text prompts and images, which has also been adopted in many image editing works [29, 44, 12]. When it comes to video editing [22, 23, 32], temporal consistency needs to be considered during attention control. Video-P2P [23] extend the cross-attention control from Prompt-to-Prompt to video editing. FateZero [32] fuses self-attention with a blending mask obtained by cross-attention features from the source prompt. However, all these works are designed for video-to-video translation editing with structure preservation. Differently, we aim for appearance-consistent video editing. Besides, none of these works explore attention control in diffusion transformers. In this paper, we are the first to analyze how the full attention in diffusion transformers could be utilized for editing, similar to U-Net diffusion models.
@@ 3.1 MM-DiT Attention Mechanism Analysis
The MM-DiT is the fundamental architecture of the current SOTA method of Text-to-image/Video models [13, 45, 14, 6], which is fundamentally distinct from prior UNet architectures since it maps text and videos into a unified sequence for attention computation. Although it has been widely utilized, the properties of its inner attention mechanism remain insufficiently explored, which restricts its applications in our multi-prompt longer video generation task. Therefore, for the first time, we conducted a comprehensive analysis of the regional attention patterns in the 3D full attention map based on the state-of-the-art video model, i.e. CogVideoX [45].
@@ 3.2 Consistent Video Generation Over Time
Figure 4: Pipeline of the proposed DiTCtrl. Our method tries to synthesize content-consistent and motion-consistent videos based on multi-prompts. The first video is synthesized with source text prompt $P_{i-1}$. During the denoising process for video synthesis, we convert the full-attention into masked-guided KV-sharing strategy to query video contents from source video $\mathcal{V}_{i-1}$, so that we can synthesize content-consistent video under the modified target prompt $P_{i}$. Note that initial latents are assumed to be 5 frames. The first three frames are used to generate the contents of $P_{i-1}$, and the last three frames are used to generate contents of $P_{i}$. The pink latent represents the overlapping frame, while the blue and green latents are used to distinguish different prompt segments.
Based on the previous analysis, we find the attention mechanism in the MM-DiT has a similar behavior as that in the UNet-like video diffusion model with our specific design. Thus, we propose the masked-guided KV-sharing strategy for consistent video generation over time for our multi-prompt video generation task.
Specifically, as shown in Fig. 4, to generate the consistent video between prompt $P_{i-1}$ and prompt $P_{i}$, we utilize the intermediate attentions from the $i-1$-th and $i$-th prompt in MM-DiT to generate the attention mask of the specific consistent object, respectively. This is achieved by averaging all the text-to-video/video-to-text parts of the 3D full attention with the given specific subject token. With these masks, we then perform mask-guided attention fusion to generate the new attention features of the prompt $P_{i}$. Inspired by MasaCtrl [9], we directly utilize the key and values from the prompt $P_{i-1}$ to guide the generation of prompt $P_{i}$ to generate the consistent appearance over time.
Formally, at step $t$, we perform a forward pass with the fixed MM-DiT backbone with prompt $P_{i-1}$ and next prompt $P_{i}$, respectively, to generate intermediate regional cross-attention maps. Then we average the attention maps across all heads and layers with the same spatial resolution $H\times W$ and temporal frames $F$. The resulting cross-attention maps are denoted as ${A^{c}_{i}}\in\mathbb{R}^{F\times H\times W\times N}$, where $N$ is the number of the textual tokens. We then obtain the averaged cross-attention map for the token correlated to the foreground object. We denote $M_{i-1}$ and $M_{i}$ as masks extracted for the foreground objects in $\mathcal{V}_{i-1}$ and $\mathcal{V}_{i}$, respectively. With these masks, we can restrict the object in $\mathcal{V}_{i}$ to query contents information only from the object region in $\mathcal{V}_{i-1}$:
@@
where $\bar{f}^{l}$ is the final attention output. Then, we replace the feature map of the current step to $f^{l}_{o}$ for further calculation.
@@ 3.3 Latent Blending Strategy for Transition
While our previous methods enable semantic consistency between clips, achieving smooth transitions between different semantic segments still needs to be carefully designed. Thus, we propose a latent blending strategy to ensure temporal coherence across different semantic segments, inspired by recent works in single-prompt long video generation [33, 46].
As illustrated in Fig. 5, our approach introduces overlapped regions between adjacent semantic video segments (video $\mathcal{V}_{i-1}$ and video $\mathcal{V}_{i}$). For each frame position in the overlapped region, we apply a position-dependent weight function that follows a symmetric distribution - frames closer to their respective segments receive higher weights while those at the boundaries receive lower weights. This weighting scheme ensures smooth transitions between different semantic contexts.
Formally, given two adjacent video segments $\mathcal{V}_{i-1}$ and $\mathcal{V}_{i}$ generated from prompts $P_{i-1}$ and $P_{i}$ respectively, we propose a latent blending strategy as follows. Let $T$ denote the number of overlapped frames between segments. For frame position $t$ in the overlapped region, we compute its blended latent features $\mathbf{z}_{t}$ as:
$\mathbf{z}_{t}=w(t)\cdot\mathbf{z}_{t,i-1}+(1-w(t))\cdot\mathbf{z}_{t,i},$ (5)
where $\mathbf{z}_{t,i-1}$ and $\mathbf{z}_{t,i}$ are latent features from $\mathcal{V}_{i-1}$ and $\mathcal{V}_{i}$ respectively, and $w(t)$ is a position-dependent triangular weight function defined as:
$w(t)=\min\left(\frac{2(t+0.5)}{T},2-\frac{2(t+0.5)}{T}\right),\quad t\in[0,T-1]$ (6)
@@
The key advantage of our approach is that it requires no additional training while effectively handling transitions between different semantic contexts. During each denoising step, we first process each segment independently, then progressively blend the latent features in the overlapped regions using position-dependent weights. This strategy maintains temporal coherence while smoothly transitioning between different semantic contexts, making it particularly suitable for multi-prompt video generation tasks.
@@ 4 Experiments
Figure 6: Generation results on given prompts by our method and baseline models. Kling is the commercial model, and Freenoise+DiT is our implementation of Freenoise on CogVideoX.
@@
We implement DiTCtrl based on CogVideoX-2B [45], which is a state-of-the-art open-source text-to-video diffusion model based on MM-DiT. In our experiments, we generate multi-prompt conditioned video, and each video clip consists of 49 frames with $480\times 720$ resolution. Moreover, we employ ChatGPT [1], to generate multiple transitions of different types. We set latent sampling frames and overlap sizes to 13 and 6 in our experiments. The experiments are conducted on a single NVIDIA A100 GPU.
4.1 Qualitative Results
We conduct comprehensive qualitative comparisons with state-of-the-art multi-prompt video generation methods [33, 40, 38] and leading commercial solutions [3]. To ensure a fair comparison, we additionally implement FreeNoise [33] on the CogVideoX backbone as an enhanced baseline.
As shown in Fig. 6, our proposed method demonstrates superior performance across three critical aspects: text-to-video alignment, temporal coherence, and motion quality. While Kling [3] exhibits impressive capabilities in high-quality video generation, it is limited to simultaneous multi-semantic blending rather than sequential semantic transitions, highlighting the significance of our work in achieving temporally evolving content across multiple prompts.
Our comparative analysis reveals distinct characteristics and limitations of existing approaches. Gen-L-Video [40] suffers from severe temporal jittering and occasional subject disappearance, compromising overall video quality. Video-Infinity [38] and FreeNoise [33] both demonstrate successful scene-level semantic changes but lack physically plausible motion - for instance, in Fig. 6, vehicles appear to be in motion while remaining spatially fixed, which is a limitation inherent to their UNet-based abilities. In contrast, FreeNoise+DiT leverages the DiT architecture’s abilities to achieve more realistic object motion but struggles with semantic transitions, resulting in noticeable discontinuities between segments. Our proposed DiTCtrl method preserves the inherent capabilities of the pre-trained DiT model while addressing these limitations, enabling smooth semantic transitions and maintaining motion coherence throughout the video sequence. For a more comprehensive evaluation, we provide additional frame-level and video-level comparisons with extensive qualitative examples in the supp.
4.2 Quantitative Results
In this section, we will first elaborate on our proposed new benchmark MPVBench for evaluating multi-prompt video generation, and then discuss the quantitative results.
@@ 4.2 Quantitative Results
Figure 7: T-SNE visualization of CLIP embeddings. Each point represents the CLIP embedding of a single video frame after dimensionality reduction. The visualization demonstrates that conventional multi-prompt videos form distinct clusters, while our method produces a more continuous distribution, indicating smoother semantic transitions.
@@
MPVBench. MPVBench contains a diverse prompt dataset and a new metric customized for multi-prompt generation. Specifically, leveraging GPT-4, we produce $130$ long-form prompts of $10$ different transition modes. Then, for multi-prompt video generation, we observe that the distribution of the CLIP features differs between single-prompt and multi-prompt scenarios. As shown in Fig. 7, the feature points of natural video follow a continuous curve, while those of two concatenated isolated videos follow two continuous curves with a breakpoint in the middle. Since the common CLIP similarity calculates the average of neighborhood similarities, the difference between natural and isolated video only occurs at the breakpoint, which becomes very small when divided by the number of frames. To address this limitation, we propose CSCV (Clip Similarity Coefficient of Variation), a metric specifically designed to evaluate the smoothness of multi-prompt transitions, defined as:
@@
where $\mathbf{x}_{i}$ denotes frame features, $\sigma$ and $\mu$ are standard deviation and average respectively. The Coefficient of Variation $CV=\sigma(s)/\mu(s)$ describes the degree of uniformity, which can largely punish the isolated situation. The function $\frac{1}{1+\lambda(\cdot)}$ projects the score to $[0,1]$, the larger the better.
Automatic Evaluation. We conduct the automatic evaluation with our MPVBench. From Table 1 one can see that our method achieves the highest CSCV score, demonstrating superior transition handling and overall stability in generation patterns. While FreeNoise ranks second with relatively strong stability, other methods significantly lag behind in this aspect, which is consistent with the T-SNE visualization of CLIP embedding as shown in Fig. 7. In terms of motion smoothness, our approach exhibits superior performance in motion quality and consistency. Regarding Text-Image Similarity metrics, although FreeNoise and Video-Infinity achieve higher scores, this can be attributed to our method’s kv-sharing mechanism, where subsequent video segments inherently learn from preceding semantic content.
As shown in Fig. 6, our design choice allows the road surface to gradually transition to snowy conditions while retaining features from the previous scene. Despite potentially lower text-image alignment scores, it ensures superior semantic continuity in the sequences. In practice, this trade-off doesn’t negatively impact the visual quality in multi-prompt scenarios, as demonstrated by our user study results presented below.
Human Evaluation. We invited 28 users to evaluate five models: Gen-L-Video [40], Video-Infinity [38], FreeNoise [33], FreeNoise+DiT and our method. We employ a Likert scale ranging from 1 (low quality) to 5 (high quality). Participants score each method considering overall preference, motion pattern, temporal consistency and text alignment over 16 videos generated by different scenarios. As clearly indicated in Table 2, generated videos from our method significantly outperform other state-of-the-art approaches in all four criteria, demonstrating superior capability in producing videos with natural semantic transitions that better align with human preferences for visual coherence and continuity. More details will be provided in the supplementary material.

### v2

@@ Abstract
Sora-like video generation models have achieved remarkable progress with a Multi-Modal Diffusion Transformer (MM-DiT) architecture. However, the current video generation models predominantly focus on single-prompt, struggling to generate coherent scenes with multiple sequential prompts that better reflect real-world dynamic scenarios. While some pioneering works have explored multi-prompt video generation, they face significant challenges including strict training data requirements, weak prompt following, and unnatural transitions. To address these problems, we propose DiTCtrl, a training-free multi-prompt video generation method under MM-DiT architectures for the first time. Our key idea is to take the multi-prompt video generation task as temporal video editing with smooth transitions. To achieve this goal, we first analyze MM-DiT’s attention mechanism, finding that the 3D full attention behaves similarly to that of the cross/self-attention blocks in the UNet-like diffusion models, enabling mask-guided precise semantic control across different prompts with attention sharing for multi-prompt video generation. Based on our careful design, the video generated by DiTCtrl achieves smooth transitions and consistent object motion given multiple sequential prompts without additional training. Besides, we also present MPVBench, a new benchmark specially designed for multi-prompt video generation to evaluate the performance of multi-prompt generation. Extensive experiments demonstrate that our method achieves state-of-the-art performance without additional training. Code is available at https://github.com/TencentARC/DiTCtrl.
Figure 1: Our method DiTCtrl takes multiple text prompts as input and demonstrates superior capability in generating longer videos with multiple events, long-range coherence and smooth transitions as output.
@@ 1 Introduction
Text-to-video (T2V) generation has made remarkable progress in the AIGC era [10, 50, 44], and breakthroughs such as Sora [30] have demonstrated impressive capabilities in generating longer videos through DiT [32] architecture and large-scale per-taining. However, feeding sequential prompts into current state-of-the-art text-to-video generation models (e.g., Kling [2], Gen3 [1], CogVideoX [48]) directly will produce isolated video sequences without natural transitions, as shown in Fig. 6 (First row). On the other hand, when multiple prompts are consolidated into a single-prompt describing long-term temporal changes, the generated results fail to capture semantic transitions effectively, as demonstrated in Fig. 6 (Second row). This limitation stems from their fundamental design and single-prompt training paradigm, making them inadequate for depicting real-world scenarios’ dynamic, multi-event nature.
@@
  • We pioneer the analysis of MM-DiT’s attention mechanism, finding that 3D full attention has similar behaviors to that of the cross/self-attention blocks in the UNet-like diffusion models, enabling mask-guided precise semantic control across different prompts for enhanced generation consistency.
@@ 2 Related Work
Training diffusion models on long videos often requires significant computational resources. Consequently, current video diffusion models are typically trained on videos with a limited number of frames. As a result, the quality of generated videos often degrades significantly during inference when generating longer videos. To address this problem, some works [18, 46, 42, 17] employ an autoregressive mechanism for long video generation. However, due to error accumulation, these methods often suffer from quality degradation after a few iterations. Alternatively, tuning-free methods [43, 36, 41, 23, 4] have been developed to extend off-the-shelf short-video diffusion models for generating long videos without additional training. For instance, Gen-L-Video [43] processes long videos as short video clips with temporal overlapping during the denoising process. FreeNoise [36] explores the influence of initial noises and conducts temporal attention fusion based on the sliding window for temporal consistency. MultiDiffusion [5] and Mimicmotion [49] introduces the latents blending strategy to achieve smooth transitions. Inspired by these works, we propose a novel KV-sharing mechanism and latent blending strategy for seamless transitions between different segments without additional training.
@@
Attention control is gaining popularity due to its ability to perform zero-shot image or video editing without the need for additional data. In the realm of image editing, MasaCtrl [9] enhances the existing self-attention mechanism in diffusion models by introducing mutual self-attention. This allows for querying correlated content and textures from source images, ensuring consistent and coherent edits. Prompt-to-Prompt [19] utilizes cross-attention layers to control the relation between text prompts and images, which has also been adopted in many image editing works [31, 47, 12]. When it comes to video editing [24, 25, 35], temporal consistency needs to be considered during attention control. Video-P2P [25] extend the cross-attention control from Prompt-to-Prompt to video editing. FateZero [35] fuses self-attention with a blending mask obtained by cross-attention features from the source prompt. However, all these works are designed for video-to-video translation editing with structure preservation. Differently, we aim for appearance-consistent video editing over time. Besides, none of these works explore attention control in diffusion transformers. In this paper, we are the first to analyze how the full attention in diffusion transformers could be utilized for video editing over time in multi-prompt video generation.
@@ 3.1 MM-DiT Attention Mechanism Analysis
The MM-DiT is the fundamental architecture of the current SOTA method of Text-to-image/Video models [13, 48, 14, 6], which is fundamentally distinct from prior UNet architectures since it maps text and videos into a unified sequence for attention computation. Although it has been widely utilized, the properties of its inner attention mechanism remain insufficiently explored, which restricts its applications in our multi-prompt longer video generation task. Therefore, for the first time, we conducted a comprehensive analysis of the regional attention patterns in the 3D full attention map based on the open-source video model, i.e. CogVideoX [48].
@@ 3.2 Consistent Video Generation Over Time
Figure 4: Pipeline of the proposed DiTCtrl. Note that initial latents are assumed to be 5 frames here. The first three frames are used to generate the contents of $P_{i-1}$, and the last three frames are used to generate contents of $P_{i}$. The pink latent represents the overlapping frame, while the blue and green latents are used to distinguish different prompt segments. Our method tries to synthesize content-consistent videos based on multi-prompts. The first video is synthesized with source text prompt $P_{i-1}$. During the denoising process for video synthesis, we convert the full-attention into masked-guided KV-sharing strategy to query video contents from source video $\mathcal{V}_{i-1}$, so that we can synthesize content-consistent video under the modified target prompt $P_{i}$.
Based on the previous analysis, we propose the masked-guided KV-sharing strategy for consistent video generation over time for our multi-prompt video generation task. As shown in Fig. 4, to generate the consistent video between prompt $P_{i-1}$ and prompt $P_{i}$, we utilize the intermediate attentions from the $i-1$-th and $i$-th prompt in MM-DiT to generate the attention masks of the specified foreground object (“knight” in Fig. 4). This is achieved by averaging the Text-Video/Video-Text parts of the 3D full attention across all heads and layers with the given object tokens, then thresholded to obtain binary masks $M$. Inspired by MasaCtrl [9], we leverage the masks to conduct mask-guided attention fusion based on the key and values from the prompt $P_{i-1}$ to generate the new attention features of the prompt $P_{i}$.
We denote $M_{i-1}$ and $M_{i}$ as masks extracted for the foreground objects in videos $\mathcal{V}_{i-1}$ and $\mathcal{V}_{i}$, respectively. With these masks, we can restrict the object in $\mathcal{V}_{i}$ to query contents information only from the object region in $\mathcal{V}_{i-1}$:
@@
where $\bar{f}^{l}$ is the final attention output. The object regions and the background regions in the current video query the content information from corresponding restricted areas rather than all the last video features.
@@ 3.3 Latent Blending Strategy for Transition
While our previous methods enable semantic consistency between adjacent video segments, achieving smooth transitions between different semantic segments still needs to be carefully designed. Thus, we propose a latent blending strategy to ensure temporal coherence across different semantic segments, inspired by recent works  [36, 49, 5].
As illustrated in Fig. 5, our approach introduces overlapped regions between adjacent semantic video segments. For each frame position in the overlapped region, we apply a position-dependent weight function that follows a symmetric distribution, in which frames closer to their respective segments receive higher weights while those at the boundaries receive lower weights. This weighting scheme ensures smooth transitions between different semantic contexts.
The blended latent feature $\mathbf{z}_{t}$ for frame $t$ is calculated as:
$\mathbf{z}_{t}=\frac{\sum_{i=1}^{n}w(t_{i})\cdot\mathbf{z}_{t_{i}}}{\sum_{i=1}^{n}w(t_{i})},$ (5)
$w(t_{i})=\min\left(\frac{2(t_{i}+0.5)}{T},2-\frac{2(t_{i}+0.5)}{T}\right),$ (6)
where $\mathbf{z}_{t_{i}}$ is latent feature corresponding to ${t}$-th frame in ${i}$-th latent segment, $n$ is number of overlapped segments, $T$ denotes the frames number of one latent segment and $w(t_{i})$ is a position-dependent weight function.
@@
To conclude, our approach employs the latent blending strategy and kv-sharing mechanism simultaneously during each denoising step. We process segment pairs sequentially, feeding them as one batch into the MM-DiT block for mask-guided kv-sharing (Fig. 4), then blend their denoised latents progressively (Fig. 5).
@@ 4 Experiments
Figure 6: Comparison of generation results across methods. Freenoise+DiT is our implementation of Freenoise on CogVideoX.
@@
Setup. We implement DiTCtrl based on CogVideoX-2B [48], which is an open-source text-to-video diffusion model based on MM-DiT. In our experimental setup, we generate multi-prompt conditioned videos, where each video clip is composed of 49 frames with a resolution of $480\times 720$. The sample step is configured to 50. The kv-sharing steps are set as [2,25], and the kv-sharing layers are specified as [25,30]. For latent sampling, the number of frames is set to 13, and the overlap size is set to 6 in our experiments. All these experiments are conducted on a single NVIDIA A100 GPU.
Baselines. We mainly compare the proposed tuning-free method to current state-of-the-art multi-prompt video generation methods [36, 43, 41] and leading commercial solutions Kling [2]. Gen-L-Video, FreeNoise, and Video-Infinity are built upon the VideoCrafter2 [11] framework. To ensure a fair comparison of base models, we implement FreeNoise [36] as an enhanced baseline by directly incorporating their noise rescheduling strategy into the CogVideoX framework.
@@
Figure 7: t-SNE visualization of CLIP embeddings. Each point represents the CLIP embedding of a single video frame after dimensionality reduction. The visualization demonstrates that conventional multi-prompt videos form distinct clusters, while our method produces a more continuous distribution, indicating smoother semantic transitions. More details and discussions will be given in the supplementary.
4.1 MPVBench
@@
MPVBench contains a diverse prompt dataset and a new metric customized for multi-prompt generation. Specifically, leveraging GPT-4, we produce $130$ long-form prompts of $10$ different transition modes. Then, for multi-prompt video generation, we observe that the distribution of the CLIP features differs between single-prompt and multi-prompt scenarios. As shown in Fig. 7, the feature points of real video (from DAVIS [34]) follow a continuous curve, while those of two concatenated isolated videos follow two continuous curves with a breakpoint in the middle. Since the common CLIP similarity calculates the average of neighborhood similarities, the difference between real video and isolated video only occurs at the breakpoint, which becomes very small when divided by the number of frames. To address this limitation, we propose CSCV (Clip Similarity Coefficient of Variation), a metric specifically designed to evaluate the transition smoothness of multi-prompt videos, defined as:
@@
where $\mathbf{x}_{i}$ denotes frame features, $\sigma$ and $\mu$ are standard deviation and average respectively. The Coefficient of Variation $CV=\sigma(s)/\mu(s)$ describes the degree of uniformity, which can largely punish the isolated situation. The function $\frac{1}{1+\lambda(\cdot)}$ projects the score to $[0,1]$, the larger the better. We also report the Text-Image similarity by using CLIP Similarity [20] to assess the alignment between given prompts and output video clips, and Motion smoothness from VBench [22] to evaluate whether the motion in the generated video is smooth, and follows the physical law of the real world.
4.2 Qualitative Results
The qualitative comparison with previous multi-prompt video generation methods is shown in Fig. 6. Notably, when multiple prompts are consolidated into a single-prompt describing long-term temporal changes, the generated result by Kling [2] fails to capture semantic transitions effectively, where the sun remains present while no clouds appear, which does not align with the text. Gen-L-Video [43] suffers from severe temporal jittering, compromising overall video quality. Video-Infinity [41] and FreeNoise [36] both demonstrate successful scene-level semantic changes but lack physically plausible motion. For instance, in Fig. 6, vehicles appear to be in motion while remaining spatially fixed, which is a limitation inherent to their UNet base-model abilities. In contrast, FreeNoise+DiT leverages the DiT architecture’s abilities to achieve more realistic object motion but struggles with semantic transitions, resulting in noticeable discontinuities between segments. Our DiTCtrl preserves the inherent capabilities of the pre-trained DiT model while addressing these limitations, enabling smooth semantic transitions and maintaining motion coherence throughout the video sequence. For a more comprehensive evaluation, we provide additional comparisons with extensive qualitative examples in the supplementary.